OpenStudy (anonymous):

Please I need help... 9. Find a linear function F(x)=ax+b for which f(2)=1 and f(-3)=11

OpenStudy (callisto):

We are given f(2)=1 and f(-3)=11 So, let's consider f(2)=1, which means when you replace x by 2, you'll get 1 from the function. So, replacing x by 2, we get f(2) = a(2) + b And we know f(2) =1. Hence, we get 1 = a(2) + b Can you use similar logic to get another equation in terms of a and b by f(-3)=11?

OpenStudy (anonymous):

Thank you :)

OpenStudy (callisto):

You're welcome. Do you understand how to do it now?